OpenStudy (anonymous):

Find an equation of a line parallel to the line 3x-2y=4 and containing the point (3,-5)

OpenStudy (anonymous):

\[2y=4\]is this wat u tryin to find

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

parallel = means the same slope re-write your ew as: 2y=3x-4 or y=3/2 x -2 so, slope your existing line is: m=3/2 so, eq of parallel through the point (3,5): y-5=3/2 *(x-3) y=3/2x -9/2 +5=3/2 x+1/2

OpenStudy (anonymous):

That makes sense.... Thank you!

OpenStudy (anonymous):

Can you help me with another one?

OpenStudy (anonymous):

shoot... I need to go soon

OpenStudy (anonymous):

Its like the one you just helped me with only perpendicular to the line 3x-2y=4 and containing points 3,-5

OpenStudy (anonymous):

try to do it yourself - I'll help. just remember that perpendicular means that your line will have slope (m2) = -1/m1... so, if re-write your eq: y=3/2 x -2 so, for perp line: m2=-2/3 eq of the line: y-(-5)=-2/3 (x-3) simplify, please

OpenStudy (anonymous):

So then y=-2/3x+6/3-5?

OpenStudy (anonymous):

y+5=-2/3 *x +2 y=-2/3 *x -3

OpenStudy (anonymous):

do you see it?

OpenStudy (anonymous):

So the equation is y=-2/3x-3?

OpenStudy (anonymous):

yep... to be exact: y=(-2x/3 ) -3

OpenStudy (anonymous):

Okay, I guess I just don't get how you suddenly got to x+2 in the end

OpenStudy (anonymous):

let's go back to: y-(-5)=-2/3 (x-3) i think that the way it written it could be confusing... let's try differently: y+5= x* (-2/3) -3* (-2/3) y+5=-2x/3 +2 (-3 in numerator & -3 in dominator are canceled out) is it better...?

OpenStudy (anonymous):

That does make better sense. Math is like a whole other language for me, lol. Thank you so much for your help

OpenStudy (anonymous):

Great! Welcome :)